Two uniform spheres, each with mass 𝑀 and radius 𝑅, touch each other. What is the magnitude of their gravitational force of attraction?Β 


1
Expert's answer
2021-12-17T18:16:43-0500

The magnitude of gravitational force of two spheres is given as follows:


"F = G\dfrac{M\cdot M}{r^2}"

where "G" is the gravitational constant, and "r" is the distance between the centers of the spheres. Since they touch each other, "r = 2R". Finally, obtain:


"F = G\dfrac{M^2}{4R^2}"